OpenStudy (amber15):

After driving 236 km, 4/5 of a trip is completed. how long is the total trip? how many kilometers must be driven to complete the trip, the entire trip is ___ km,and how many ___km , remaining

OpenStudy (anonymous):

236 = (4/5)*x x is the length of the trip. x = 295 km total. amount remaining : 59 km

OpenStudy (anonymous):

Let the total distance to be covered be = x \[\frac{ 4 }{ 5 } \times x = 236\] Therefore, x = 295 And the number of km remaining = 295 - 236 = 59km. :-)

OpenStudy (amber15):

Thank you both